OpenStudy (anonymous):

How to find the maximum value of g(x) = 2 cos(2x - π) + 4?

OpenStudy (anonymous):

what is the maximum value of cos?

OpenStudy (anonymous):

isn't one?

OpenStudy (anonymous):

so the max value is ¨6

OpenStudy (anonymous):

@RaphaelFilgueiras can you explain more, how to know that it's 6? :o I forgot this.

OpenStudy (anonymous):

-1<=cos(u)<=+1, then the max value of g(x), is reached when cos(2x-pi)=1

OpenStudy (anonymous):

ohh thanks xD @RaphaelFilgueiras

OpenStudy (anonymous):

I was thinking in terms of f(x) = a (cosb(x-h)) + k max= a+k

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

you can differentiate it too

OpenStudy (anonymous):

but idk if it's always true, thanks for proving it thought! <3

OpenStudy (anonymous):

wait so how do y'all know it's 6? @study100 @RaphaelFilgueiras

OpenStudy (anonymous):

tried googling it and it said if a < 0 then the max is k, but it didn't explain how to find the max value when a > 0

OpenStudy (anonymous):

http://web.psjaisd.us/auston.cron/ABCronPortal/GeoGebraMenu/GeogebraFiles/trigonometry/graphics/cosineDefn.PNG refer to image, you see that a+k = max point see the green and purple line add up to max

OpenStudy (anonymous):

a= 2 k= 4 so 2+4 = 6 but raphael has a really good way to prove that too.

OpenStudy (anonymous):

so this method would only be used if a > 0? @study100

OpenStudy (anonymous):

This is the other equation I needed to find the max value of. f(x) = -4(x - 6)^2 + 3 . I said it was 3, is that correct? Or would it be -1?

OpenStudy (anonymous):

oh when a is less than 0 all you need to do is forget about the negative sign in a. maximum of ALL COS FUNCTION = absolute value of a plus c = l a I + k

OpenStudy (anonymous):

I assume you know absolute value?

OpenStudy (anonymous):

plus k***

OpenStudy (anonymous):

Proof, here I graphed -acos... for you

OpenStudy (anonymous):

still 6, see? :DD

OpenStudy (anonymous):

y = a(x-h) ^2 + k max is vertex vertex = (h,k) h= 6 k= 3 So the exact point is at (6,3) for max

OpenStudy (anonymous):

Ahhh I'm so confused lol

OpenStudy (anonymous):

first or second?

OpenStudy (anonymous):

That's why I thought when a < 0, the max value is automatically k

OpenStudy (anonymous):

in the second one the first therm is always negative then the max is when it's zero, so the max is 3

OpenStudy (anonymous):

Method 2

OpenStudy (anonymous):

So for this equation the max value WOULD be 3, right? f(x) = -4(x - 6)^2 + 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And so the max value of g(x) = 2 cos(2x - π) + 4? is 12?

OpenStudy (anonymous):



OpenStudy (anonymous):

if you use the y = c- (b^2/2a) it will generate the same answer